Problem 15-11
WACC and Optimal Capital Structure

F. Pierce Products Inc. is considering changing its capital structure. F. Pierce currently has no debt and no preferred stock, but would like to add some debt to take advantage of low interest rates and the tax shield. Its investment banker has indicated that the pre-tax cost of debt under various possible capital structures would be as follows:

Market Debt-
to-Value
Ratio (wd)
Market Equity-to-Value
Ratio (ws)
Market Debt-
to-Equity
Ratio (D/S)
Before-Tax Cost of Debt (rd)
0.0 1.0 0.00 7.0%
0.2 0.8 0.25 8.0  
0.4 0.6 0.67 10.0  
0.6 0.4 1.50 12.0  
0.8 0.2 4.00 15.0  

F. Pierce uses the CAPM to estimate its cost of common equity, rs. The company estimates that the risk-free rate is 7%, the market risk premium is 6%, and the company's tax rate is 30%. F. Pierce estimates that its beta now (which is "unlevered" since it currently has no debt) is 1. Based on this information, what is the firm's optimal capital structure, and what would the weighted average cost of capital be at the optimal capital structure? Do not round intermediate calculations. Round your answers to two decimal places.

DEBT %
EQUITY %
WACC %
Answer #1

A B C D G H 1 D/Cap E/Cap D/E Bond Rating Before tax cost of debt Levered beta Cost of equity WACC 2 0 1 0 7% 1 13.00000% 13.

H A B C 1 D/Cap E/Cap D/E 2 0 =1-A2 =A2/B2 3 0.2 =1-A3 =A3/B3 4 0.4 =1-A4 0.67 5 0.6 =1-A5 =A5/B5 6 0.8 =1-A6 =A6/B6 D Bond R

Debt 40%
Equity 60%
WACC 12.288%